OpenStudy (anonymous):

An unopened bag of candy weighed 5 1/2 lb. After opening the bag and eating some of the candy, the bag weighed 2 1/4 lb. How many pounds of candy were eaten?

OpenStudy (anonymous):

We need to subtract fractions

OpenStudy (anonymous):

okay

OpenStudy (haleyelizabeth2017):

what is 0.5-0.25? (This is the easier way to subtract the fractions.

OpenStudy (anonymous):

so it would be 3 1/4?

OpenStudy (haleyelizabeth2017):

Yes!

OpenStudy (anonymous):

\[\frac { 1 } { 2 }{ 5 } - \frac { 1 } { 4 }{ 2 } = 3 2/8 = 1/4\]

OpenStudy (anonymous):

the answer is 3 1/4

OpenStudy (anonymous):

thank you guys :D

OpenStudy (haleyelizabeth2017):

No problem!